# The Peekaboo galaxy: new SALT spectroscopy and implications of archive HST data

### v1

@@ 4 Discussion
4.1 Metallicity of Peekaboo in the context
@@
Summarising the new determination of O/H in the Peekaboo dwarf, which appeared close to the earlier value, but with much improved accuracy, 12+log(O/H)dir = 6.99 0.06 dex, we compare this value to the other known dwarfs with the lowest metallicities in the LV. The best to our knowledge, within the LV (R11 Mpc), apart Peekaboo, there were found 10 XMP dwarfs with 12+log(O/H) 7.17 dex (or Z(gas) 0.03 Z⊙). Six of them are published in Pustilnik et al. (2010, 2021, 2024); Skillman et al. (2013); Izotov et al. (2012). See Table 6 in Appendix, where we compile their properties. Five of these six dwarfs reside within the nearby voids from Pustilnik et al. (2019). In the framework of the ongoing project ”Studying the void galaxies in the Local Volume”, we found 4 new LV XMP dwarfs, currently prepared for publication. Only four of the total 11 LV XMP galaxies (including Peekaboo) have their O/H derived with the direct method.
@@
One dwarf, Leo P, with 12+log(O/H) =7.17 dex (Skillman et al. 2013) is the nearest known XMP galaxy, at 1.6 Mpc from the Local Group center. Its proximity allowed one to study its massive hot XMP star O7-8 V with the ground-based spectroscopy and to obtain the important constraints of its wind properties (Telford et al. 2023, 2024). The remaining Local Volume XMP dwarfs are situated much farther, at the distances of 7–11 Mpc. Thus, spectroscopy of their luminous stars should be available with HST and upcoming ELTs.
4.2 Hot and luminous stars. Blue outliers
@@
The proximity of Peekaboo and the high resolution of the HST images allow one to identify the hottest massive stars, including stars ionising the E and W H ii regions. We separate in the diagram MV versus (Fig. 8), a region, where all O-stars (for luminosity classes from V to I) fall according to the state-of-art models at metallicity of Z = Z⊙/10, presented by Lorenzo et al. (2025). This is a rectangular with boundaries of from –0.32 to –0.25 mag, and MV between –3.30 and –6.15 mag. The list of these stars with their additional parameters is presented in Appendix. In Fig. 8, we also identify four supergiant stars, the evolved massive stars, conditionally with M –6.0 mag (or 21.4 23.5 mag). They probably represent the upper part of Blue Helium Burning (BHeB) stars according to McQuinn et al. (2012). They are interesting for detailed studies as the nearest evolved massive stars with the record-low metallicity.
@@
In particular, they can be observed spectroscopically at the superb conditions (narrow slit and seeing of 0.5 arcsec) with the Very Large Telescope at ESO. In addition, there is a good opportunity to check their possible large-amplitude optical variability via monitoring of the entire galaxy and analysing light variations of the small regions around these supergiants. A successful example of monitoring of variability of the XMP LBV and other supergiants in H ii regions of the XMP galaxy DDO68 (12+log(O/H) = 6.98 – 7.3 dex, Pustilnik et al 2005; Izotov & Thuan 2007; Annibali et al. 2019) was recently presented in Pustilnik et al. (2024); Pustilnik, Perepelitsyna (2025).
@@
Another type of evolved massive stars can show up in versus diagram as ’blue outliers’ (e.g. Lorenzo et al. 2022). These stars with M –3.5 to –5 mag and colours bluer than the blue edge for the Main Sequence stars (namely, of –0.32, Lorenzo et al. (2025)) can be the Helium-burning WO stars with the surface temperatures up to of 150–200 kK (Lorenzo et al. 2022, and references therein).
4.3 Properties of Peekaboo H ii regions
@@
In Appendix section B we present the distribution of H flux along the slit, which looks like two close ’gaussians’ with FWHM of 3.4 and 3.7 arcsec. Their sizes along the slit (corrected for seeing of 1.5 arcsec) are FWHM 3.05 and 3.4 arcsec, or 116 pc and 128 pc. In the colour HST image one does not see such clear extended structures. These sizes probably represent the extent of the (smoothed due to seeing) collection of several smaller H ii regions, elongated along the galaxy major axis. Several of the brightest smaller H ii regions are marked in Fig. 1 and their parameters are summarised in Table 5.
@@
As the two-component velocity profiles of all emission lines in the E H ii region evidence, this cell of ionised gas indeed has the complex origin. Roughly 30% of its nebular emission is related to the ionised gas, moving with the radial velocity of about –60 km s-1 (’blue’ component hereafter) relative to the main body, with the system velocity of V 716 km s-1 (Karachentsev et al. 2023). Both the radial velocities, of the E region more massive ’red’ velocity component, and that of the W H ii region are close to the system one.
@@
This ionised gas should be connected with a fast-moving hot star. Such massive stars, products of co-evolution in binary systems are well known in the Galaxy. Due to the ’poor’ spatial resolution of the long-slit spectra, one can not directly address the size of the ’blue’ component. However, since the ’blue’ component gas moves together with its exciting massive star, one can suggest that this shell is related to the mass loss from the fast-moving star, occurred after the epoch when it was thrown out from a close binary (or a cluster). The abundance pattern in E(blue), similar to those in the other regions, indicates that its nebular emission originates in the swept-up ambient gas shell. The case of the complex nature of the E H ii region partly resembles the case of the H ii region A2 in IC4662. In this region, the echelle spectroscopy revealed three components, of which one, related to a known WR star, moves with the relative velocity of V 20 km s-1 (Kniazev 2025).
@@
In Appendix (Sect. B) we estimate luminosities of the line H of the three emission-line subsystems. They are (7.5, 8.4 and 14.8)1036 erg s-1 for the W, E(blue) and E(red) subsystems, respectively. These luminosities can be used to estimate the conditional number of hot stars, ionising these regions. For further calculations, we adopt the relation between L(H) and the related flux of ionising quanta Q0: L(H) = 4.7610-13 Q0. Resulting values of Q0 for the three subsystem are as follows: (1.6, 1.8 and 3.2) 1049 phot s-1. Recently Telford et al. (2023) estimated the value of Q0 for the lowest accessible metallicity O7V star LP26 in Leo P (12+log(O/H)=7.17 dex). This number (3.740.67)1048 phot s-1, is 2.5 times lower than the value, usually adopted for O7V star with Z = Z⊙. That is, we need from 3 to 6 O7V XMP stars per a subsystem to provide the required amount of ionising photons.
@@
The comparison of star positions (marked with their numbers) in the HST image (top panel of Fig. 1) and their positions in the diagram of Fig. 8, reveals within the borders of the E region only one certain luminous O star (No. 7) and about 1-mag fainter probable O-stars No. 5 and 6. On the other hand, two likely luminous WR stars (No. 2 and 4) reside in the middle of the E region. One more, candidate very hot WR star No. 1 is close to No. 2. In particular, their extremely blue colours are consistent with those of WO stars, for which their temperatures are estimated to be in the range of 150-210 kK (e.g. Tramper et al. 2015). Their parameter Q0 falls in the range of (0.8–3)1049 phot s-1. That is, even one such WO star is sufficient to provide ionising photons for each of the three H ii subsystems.
@@
Since observations indicate that most of WR stars are situated on the edges of the related H ii regions (e.g., Crowther & Bibby 2009), this implies that they were ejected from the parent cluster. Therefore, one can expect that the ’fast-moving’ ’blue’ subsystem of E region is related with one of the above WR stars. The spectra of this region with a seeing of 0.5 arcsec could help to disentangle various subsystems within this region.
@@
Apart these two elongated E and W complexes, numerous small blue non-stellar objects (’nebulosities’) are distributed along the whole dwarf body. Their typical sizes of 0.2–0.5 arcsec correspond to radii of 7–15 pc, or to Strömgren radii of 4–8 pc. Such small H ii regions are probably ionised and excited by individual hot massive stars. Their () colours are bluer than those of hot O-type stars. Such very blue () colours are predicted for the nebular emission in low-metallicity high-temperature H ii regions in Zackrisson et al. (2001); Anders & Fritze-v. Alvensleben (2003). In particular, Anders & Fritze-v. Alvensleben (2003) show that colour () for young starbursts with ZZ⊙/50 can vary in the range of –0.7 to –0.25 mag for ages from 5 to 15 Myr due to the very strong contribution of emission lines in -band. This is exactly the range which we see for the five regions N1–N5, shown in Fig. 8 and in Tab. 5. As the lowest metallicity individual H ii regions in the nearby Universe, they deserve spectroscopy with HST and/or ELTs.
@@ 5 Conclusions
Summarising presented above new data on the dwarf galaxy Peekaboo and their discussion, we arrive at the following conclusions:
@@
The new spectral data for the two H ii regions of Peekaboo dwarf allow us to uncover three ionised gas systems. Two systems, the E and W, are separated spatially, at 4 arcsec between their centres (140 pc), approximately along the major axis. In the E H ii region we see two ’co-spatial’ kinematically separated gas systems with the radial velocities of Vhel of 720 km s-1 and 654 km s-1(’red’ and ’blue’ components).
@@
Thanks to the improved S-to-N ratio in the principal faint line [Oiii]4363Å of the E region (red), we derive its O/H with the direct (Te) method: 12+log(O/H) = 6.990.06 dex. For the E(blue) subsystem and for the W H ii region, the line [Oiii]4363Å is undetected. Their O/H, derived via the empirical ”strong-line” method, appear very close to that of the E(red) subsystem with (O/H)(Te). Pekaboo galaxy is the lowest metallicity dwarf within the Local Volume and its environs, among galaxies with gas O/H derived via the direct method.
@@
We derive also abundances of the other heavy elements X (Ne, Ar, N, S), which are consistent with the average X/O, typical of galaxies in the lowest metallicity range.
@@
Using the EDD coordinates, magnitudes and colours of individual stars in the body of Peekaboo dwarf, we identify the hot massive stars (including candidate WO stars), tentatively ionising its H ii regions, as well as four supergiants in this galaxy. These supergiants, as the brightest and the lowest-metallicity evolved massive stars, are the valuable targets for the detailed studies with HST and the upcoming ELTs as well as for the ground-based monitoring of their possible large-amplitude variability. They can be good probes to confront them with the modern models of stellar evolution at Z 0.02 Z⊙.
@@ Acknowledgements.
Appendix A HST image of Peekaboo dwarf with the slit position
In the top panel of Fig. 1 we show the zoomed F606W HST image of Peekaboo dwarf. The position of the 1.5-arcsec-width long slit is superimposed. We also overplot approximate boundaries of W and E H ii regions with arcs, defined through the distance from the ’reference’ star. The hottest and supergiant stars are marked to show their location relative to W and E H ii regions.
@@ Appendix A HST image of Peekaboo dwarf with the slit position
Appendix B 2D spectra, sizes of H ii regions and related issues
@@
The fluxes of emission H line of W and E regions are 8.5 and 26.3, respectively (in units of 10-16 erg cm-2 s-1). The line flux for the E region comprises of two components, with the ’blue’ component part of 9.6 and for the ’red’ component of 16.7. The related line luminosities, after correction for the measured extinction in the H ii regions, C(H)0.18 dex, or factor of 1.6, are as follows: for W region: 7.51036, for E region: 23.21036 erg s-1, with division of (8.4 and 14.8)1036 for blue and red components, respectively.
Appendix C Extracted 1d spectra of E and W H ii regions
In Figure 4 we show 1d spectra of Peekaboo E and W regions with grism PG900, in which we mark the main emission lines used for the abundance determination (Hydrogen Balmer lines and lines of O, Ne, N, S, Ar ions). The spectra are flux-calibrated via the known Legacy DR10 photometry of the reference star. In Figure 5, a similar 1d spectrum is shown, with the higher spectral resolution with the use of grism PG3000. The faint auroral line [Oiii]4363 Å is well seen near H in the spectrum of the eastern H ii region.
Appendix D Line intensities and element abundances in the E and W H ii regions of Peekaboo dwarf

### v2

@@ 4 Discussion
4.1 Metallicity of Peekaboo in context
@@
The new determination of O/H in the Peekaboo dwarf is close to the earlier value but has a much improved accuracy: 12+log(O/H)dir = 6.99 0.06 dex. We compared this value with that of known dwarfs that have the lowest metallicities in the LV. To the best of our knowledge, within the LV (R11 Mpc), apart from Peekaboo, ten XMP dwarfs with 12+log(O/H) 7.17 dex (or Z(gas) 0.03 Z⊙) have been found. Six of them are published in Pustilnik et al. (2010, 2021, 2024); Skillman et al. (2013); Izotov et al. (2012). (See Table 6 in Appendix, where we compile their properties.) Five of these six dwarfs reside within the nearby voids from Pustilnik et al. (2019). In the framework of the ongoing project ’Studying the void galaxies in the Local Volume’, we found four new LV XMP dwarfs, currently prepared for publication. Only four of the total 11 LV XMP galaxies (including Peekaboo) have their O/H derived with the direct method.
@@
The dwarf Leo P, with 12+log(O/H) =7.17 dex (Skillman et al. 2013), is the nearest known XMP galaxy, at 1.6 Mpc from the Local Group centre. Its proximity allows for study of its massive hot XMP star O7-8 V with ground-based spectroscopy and important constraints of its wind properties to be obtained (Telford et al. 2023, 2024). The remaining LV XMP dwarfs are situated much farther away, at the distances of 7–11 Mpc. Thus, spectroscopy of their luminous stars should be available with HST and upcoming ELTs.
4.2 Hot and luminous stars, and blue outliers
@@
The proximity of Peekaboo and the high resolution of the HST images allow one to identify the hottest massive stars, including stars ionising the E and W H ii regions. We separate in the diagram of MV versus in Fig. 8 a region where all O-stars (for luminosity classes from V to I) fall according to the state-of-the-art models at a metallicity of Z = Z⊙/10 presented by Lorenzo et al. (2025). This area is indicated by a rectangle with boundaries of from –0.32 to –0.25 mag and MV between –3.30 and –6.15 mag. A list of the stars in this area and their additional parameters is presented in the appendix. In Fig. 8, we also identify four supergiant stars, the evolved massive stars, conditionally with M –6.0 mag (or 21.4 23.5 mag). They probably represent the upper part of blue helium burning stars, according to McQuinn et al. (2012). They are interesting for detailed studies, as they represent the nearest evolved massive stars with a record-low metallicity.
@@
In particular, they can be observed spectroscopically at superb conditions (i.e. narrow slit and a seeing of 0.5 arcsec) with the Very Large Telescope at ESO. In addition, there is a good opportunity to check their possible large-amplitude optical variability via monitoring of the entire galaxy and to analyse the light variations of the small regions around the supergiants. A successful example of monitoring the variability of the XMP LBV and other supergiants in H ii regions of the XMP galaxy DDO68 (12+log(O/H) = 6.98 – 7.3 dex, Pustilnik et al 2005; Izotov & Thuan 2007; Annibali et al. 2019) was recently presented in Pustilnik et al. (2024); Pustilnik, Perepelitsyna (2025).
@@
Another type of evolved massive star can show up in a versus diagram as a ’blue outlier’ (e.g. Lorenzo et al. 2022). Such a star has M –3.5 to –5 mag and colours bluer than the blue edge of the Main Sequence stars (namely, –0.32; Lorenzo et al. (2025)), and they can be helium burning WO stars with a surface temperature of up to 150–200 kK (Lorenzo et al. 2022, and references therein).
4.3 Properties of Peekaboo’s H ii regions
@@
In section B of the appendix, we present the distribution of H flux along the slit, which looks similar to two close Gaussians with FWHMs of 3.4 and 3.7 arcsec, respectively. Their sizes along the slit (corrected for a seeing of 1.5 arcsec) are FWHM 3.05 and 3.4 arcsec, or 116 pc and 128 pc. In the colour HST image, one does not see such clear extended structures. These sizes probably represent the extent of the (smoothed due to seeing) collection of several smaller H ii regions elongated along the galaxy major axis. Several of the smaller brightest H ii regions are marked in Fig. 1, and their parameters are summarised in Table 5.
@@
As the two-component velocity profiles of all emission lines in the E H ii region evidence, this cell of ionised gas indeed has a complex origin. Roughly 30% of its nebular emission is related to the ionised gas, moving with a radial velocity of about –60 km s-1 (’blue’ component hereafter) relative to the main body, which has a system velocity of V 716 km s-1 (Karachentsev et al. 2023). The radial velocities of the E region’s more massive ’red’ velocity component and that of the W H ii region are close to the system one.
@@
The ionised gas of the ’blue’ component should be connected with a fast-moving hot star. Such massive stars, which are products of co-evolution in binary systems, are well known in the Galaxy. Due to the ’poor’ spatial resolution of the long-slit spectra, the size of the ’blue’ component cannot be addressed directly. However, since the ’blue’ component gas moves together with its exciting massive star, one can suggest that this shell is related to the mass loss from the fast-moving star that occurred after the epoch when it was thrown out from a close binary (or a cluster). The abundance pattern in E(blue), similar to those in the other regions, indicates that its nebular emission originates in the swept-up ambient gas shell. The case of the complex nature of the E H ii region partly resembles the case of the H ii region A2 in IC4662. In this region, the echelle spectroscopy revealed three components, and one of them is related to a known WR star and moves with a relative velocity of V 20 km s-1 (Kniazev 2025).
@@
In the appendix (Sect. B), we estimate the luminosities of the line H of the three emission-line subsystems. Respectively, they are 7.5, 8.4, and 14.81036 erg s-1 for the W, E(blue), and E(red) subsystems. These luminosities can be used to estimate the conditional number of hot stars ionising these regions. For further calculations, we adopted the relation between L(H) and the related flux of ionising quanta Q0: L(H) = 4.7610-13 Q0. The resulting values of Q0 for the three subsystems are as follows: (1.6, 1.8, and 3.2) 1049 phot s-1. Recently, Telford et al. (2023) estimated the value of Q0 for the lowest accessible metallicity O7V star LP26 in Leo P (12+log(O/H)=7.17 dex). This number, (3.740.67)1048 phot s-1, is 2.5 times lower than the value usually adopted for the O7V star with Z = Z⊙. That is, we need from three to six O7V XMP stars per subsystem to provide the required amount of ionising photons.
@@
Comparison of the star positions (marked with their numbers) in the HST image (top panel of Fig. 1) and their positions in the diagram of Fig. 8 revealed within the borders of the E region only one certain luminous O star (No. 7) and about 1-mag fainter probable O stars (No. 5 and No. 6). Besides, two likely luminous WR stars (No. 2 and No. 4) reside in the middle of the E region. One candidate very hot WR star, No. 1, is close to No. 2. In particular, their extremely blue colours are consistent with those of WO stars, for which their temperatures are estimated to be in the range 150-210 kK (e.g. Tramper et al. 2015). Their parameter Q0 falls in the range of (0.8–3)1049 phot s-1. That is, even one such WO star is sufficient to provide ionising photons for each of the three H ii subsystems.
@@
Since observations indicate that most WR stars are situated on the edges of the related H ii regions (e.g. Crowther & Bibby 2009), this implies that they were ejected from their parent cluster. Therefore, one can expect that the ’fast-moving’ ’blue’ subsystem of the E region is related to one of the above WR stars. The spectra of this region with a seeing of 0.5 arcsec could help disentangle various subsystems within this region.
@@
Apart from these two elongated E and W complexes, numerous small blue non-stellar objects (’nebulosities’) are distributed along the whole dwarf body. Their typical sizes of 0.2–0.5 arcsec correspond to linear sizes of 7–15 pc, or to Strömgren radii of 4–8 pc. Such small H ii regions are probably ionised and excited by individual hot massive stars. Their () colours are bluer than those of hot O-type stars. Such very blue () colours were predicted for the nebular emission in low-metallicity high-temperature H ii regions in Zackrisson et al. (2001); Anders & Fritze-v. Alvensleben (2003). In particular, Anders & Fritze-v. Alvensleben (2003) showed that the colour () for young starbursts with ZZ⊙/50 can vary in the range of –0.7 to –0.25 mag for ages from 5 to 15 Myr due to the very strong contribution of emission lines in the -band. This is exactly the range we observed for the five regions N1–N5 shown in Fig. 8 and in Tab. 5. As the lowest metallicity individual H ii regions in the nearby Universe, these objects deserve spectroscopy with HST and/or ELTs.
@@ 5 Conclusions
The goals of this work was to improve the accuracy of Peekaboo’s gas metallicity and to address properties of its massive stars. Based on the new data, we have presented on the dwarf galaxy Peekaboo and their discussion, we arrived at the following conclusions:
@@
The new spectral data for the two H ii regions of the Peekaboo dwarf allowed us to uncover three ionised gas systems. Two systems, E and W, are separated spatially by 4 arcsec between their centres (140 pc) approximately along the galaxy major axis. In the E H ii region, we observed two ’co-spatial’ kinematically separated gas systems with radial velocities of Vhel of 720 km s-1 and 654 km s-1(’red’ and ’blue’ components).
@@
Thanks to the improved S-to-N ratio in the principal faint line [Oiii]4363Å of the E region (red), we derived its O/H with the direct (Te) method: 12+log(O/H) = 6.990.06 dex. For the E(blue) subsystem and for the W H ii region, the line [Oiii]4363Å is undetected. Their O/H values, derived via the empirical ’strong-line’ method, appear very close to that of the E(red) subsystem with (O/H)(Te). The Peekaboo galaxy is the lowest metallicity dwarf within the LV and its environs among galaxies with gas O/H derived via the direct method.
@@
We also derived the abundances of other heavy elements X (Ne, Ar, N, S), and they are consistent with the average X/O typical of galaxies in the lowest metallicity range.
@@
Using the EDD coordinates, magnitudes, and colours of individual stars in the body of the Peekaboo dwarf, we identified the hot massive stars (including candidate WO stars) tentatively ionising its H ii regions as well as four supergiants in this galaxy. These supergiants, as the brightest and the lowest-metallicity evolved massive stars, are valuable targets for detailed studies with HST and the upcoming ELTs as well as for ground-based monitoring of their possible large-amplitude variability. Their colours and spectral features as well as their variability could be suitable data to confront them with predictions of the modern models of stellar evolution at Z 0.02 Z⊙.
Thus, the Peekaboo dwarf is among the most intriguing dwarf galaxies in the Local Volume, and as such, it warrants intensive, multi-method, in-depth study. Its properties – ranging from high-resolution Hi gas morphology and kinematics, through compact star-forming regions, to the lowest-metallicity known massive stars revealed via resolved space-based photometry and spectroscopy – are expected to significantly advance our understanding of the first building blocks of the early Universe.
@@ Acknowledgements.
Appendix A Hubble Space Telescope image of the Peekaboo dwarf with the slit position
In the top panel of Fig. 1 we show the zoomed-in F606W HST image of the Peekaboo dwarf. The position of the 1.5-arcsec-width long slit is superimposed. We also overplot approximate boundaries of W and E H ii regions with arcs, defined through the distance from the ’reference’ star. The hottest and supergiant stars are marked to show their location relative to W and E H ii regions.
@@ Appendix A Hubble Space Telescope image of the Peekaboo dwarf with the slit position
Appendix B Two-dimensional spectra, sizes of H ii regions, and related issues
@@
The fluxes of emission H line of W and E regions are 8.5 and 26.3, respectively (in units of 10-16 erg cm-2 s-1). The line flux for the E region comprises two components, with the ’blue’ component part of 9.6 and the ’red’ component of 16.7. The related line luminosities, after correction for the measured extinction in the H ii regions, C(H)0.18 dex, or factor of 1.6, are as follows. For the W region, it is 7.51036, and for the E region, it is 23.21036 erg s-1, with the division of (8.4 and 14.8)1036 for the blue and red components, respectively.
Appendix C Extracted one-dimensional spectra of the E and W H ii regions
In Figure 4 we show 1D spectra of Peekaboo’s E and W regions with grism PG900, in which we mark the main emission lines used for the abundance determination (Hydrogen Balmer lines and lines of O, Ne, N, S, Ar ions). The spectra are flux-calibrated via the known Legacy DR10 photometry of the reference star. In Figure 5, a similar 1D spectrum is shown, with the higher spectral resolution with the use of grism PG3000. The faint auroral line [Oiii]4363 Å is well seen near H in the spectrum of the eastern H ii region.
Appendix D Line intensities and element abundances in the E and W H ii regions of the Peekaboo dwarf